OpenStudy (anonymous):

Solve the following system of equations using any method 3x+14y=−78 12x+12y=84

OpenStudy (anonymous):

What do you think the solution is?

OpenStudy (anonymous):

i have no clue, can you help?

OpenStudy (anonymous):

Have you done simultaneous equations before?

OpenStudy (anonymous):

i dont think so

OpenStudy (anonymous):

Have you solved simple equations like 4x = 8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You'll be fine then.

OpenStudy (jdoe0001):

well... if you haven't covered "system of equations", how are you expected to solve it then?

OpenStudy (jdoe0001):

or using what "method" are you expected to solve it, that is, what chapter is this covering?

OpenStudy (anonymous):

What you need to do is simple, you just rearrange one of the equations so you have y or x on one side of the equation. Can you do that?

OpenStudy (anonymous):

Then I can take you through the next step after that.

OpenStudy (anonymous):

Lets solve for y first 3x+14y=−78 12x+12y=84 Times the top equation by -4 3x+14y=−78 -4(3x+14y)=−78 *-4 -12x - 56y= 312 Now you have a new equation and you can use the cancellation method to solve. Solve for y. Can you finish? -12x - 56y = 312 12x + 12y = 84

OpenStudy (jdoe0001):

if you haven't covered the material, I'd think this exercise doesn't apply to you then and your teacher should know better than to assign material he/she hasn't covered so... you may want to do that first

OpenStudy (jdoe0001):

can we do it for you? sure, do we know how to? yes does that help you? nope

OpenStudy (anonymous):

@lindsey_quick are you still with me?

OpenStudy (anonymous):

We have this now. -12x - 56y = 312 12x + 12y = 84 To solve for y we can use the cancellation method. -12x and 12x will cancel out because -12x + 12x = 0x = 0 so now we have the following equations -56y = 312 12y = 84 Can you finish solving for y ???

OpenStudy (anonymous):

i got 5.57 and 7 ?

OpenStudy (anonymous):

-56y = 312 12y = 84 What does -56y + 12y = ?

OpenStudy (anonymous):

-44y

OpenStudy (anonymous):

Correct!! Now what does 312 + 84 = ?

OpenStudy (anonymous):

396

OpenStudy (anonymous):

Correct!!! So -56y = 312 12y = 84 -56y + 12y = -44 and 312 + 84 = 396 so we have now -44y = 396 Now we need to get y by its self. since -44y is multiplication, we divide each side by - 44 \( \huge \frac{-44y}{-44} = \frac{396}{-44} \) \( \huge y = \frac{396}{-44} \) What do you get when you divide 396 by -44?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

Close, you forgot your - symbol but yes -9. Ok we found our y value to the solution now to find the x value just put the y value into one of the original equation and solve for x. I am going to use 3x+14y=−78 3x+14(-9)=−78 Can you solve for x or do you still need a little help?

OpenStudy (anonymous):

i think i got it from here , thanks

OpenStudy (anonymous):

3x+14(-9)=−78 3x -126 = - 78 3x -126 + 126 = - 78 + 126 3x = -78 + 126 3x = 48 \( \huge \frac{3x}{3} = \frac{48}{3} \) \( \huge x = \frac{48}{3} \) \( \huge x = 16 \) So the solution is ( 16, -9) Good job! Practice makes perfect. Keep working at them and you will be a pro at them!!!